# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v1

@@ V Solving SOCP-Reformulation of MVIE via Affine Scaling Algorithm
In this section, we propose a practically efficient algorithm for MVIE in a low dimension but with a large number of constraint. By directly optimizing the Cholesky factorization of the ellipsoid matrix, we reformulate the MVIE problem into a Second-Order Conic Programming (SOCP) form. Then we provide the solution of using Affine Scaling algorithm for solving SOCP, achieving efficient solving of MVIE.
@@
where $\operatorname*{diag}({\cdot})$ indicates either constructing a diagonal matrix or taking all diagonal entries of a square matrix, and $A_{\mathcal{P}}$, $b_{\mathcal{P}}$ are the coefficients of the halfspace constituting the convex polyhedra $\mathcal{P}$. However, this program still imposes orthonormality constraints on $A_{\mathcal{E}}$. Additionally, if we use $A_{\mathcal{E}}D_{\mathcal{E}}$ as the decision variable, (V) becom an SDP form which is often used to solve MVIE [1, 16, 19].
Noting that $A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}$ is always positive definite for the optimal solution of the non-degenerate problem (V). Thus the Cholesky factorization $A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}=L_{\mathcal{E}}L_{\mathcal{E}}^{\mathrm{T}}$ is unique [37] for this solution, where $L_{\mathcal{E}}$ is a lower triangular matrix with positive diagonal entries. If we treat $L_{\mathcal{E}}$ as decision variables, since $A_{\mathcal{E}}$ is orthonormal, the objective (30a) and constraints (30b) can be written as
$\det({D_{\mathcal{E}}})=\sqrt{\det({A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}})}=\det({L_{\mathcal{E}}}).$ (31)
Consequently, the orthonormality constraint (30d) on $A_{\mathcal{E}}$ is avoided, and (V) is equivalent to
@@
$~{}L_{\mathcal{E}}~{}\mathrm{is~{}lower~{}triangular},$ (32c)
Now the objective $\det({L_{\mathcal{E}}})$ is simply the product of the diagonal entries. Denoting the hypograph of geometric mean by
@@
thus maximizing the volume of the ellpsoid $\mathcal{E}$ is equivalent to maximizing $t$ with the constraint $({\operatorname*{diag}({L_{\mathcal{E}}}),t})\in\mathcal{K}_{1/n}$. As for the constraints (30b), we denote the SOC as $\mathcal{K}_{n}$ and describe the Cartesian product of $m$ SOC as $\mathcal{K}_{n}^{m}$,
@@
The MVIE problem now becomes
@@
$~{}L_{\mathcal{E}}~{}\mathrm{is~{}lower~{}triangular}.$ (36d)
$\mathcal{K}_{1/n}$ can also be represented by SOC using additional $O(n)$ variables and cones of $\mathcal{K}_{3}$ [16].
Thus, we reformulate MVIE into a pure SOCP from (V). For brevity, we denote the it as
@@
where $x\in\mathbb{R}^{\bar{n}}$, $c_{\mathcal{K}}\in\mathbb{R}^{\bar{n}}$, ${c}_{i}\in\mathbb{R}^{\bar{n}}$, ${d}_{i}\in\mathbb{R}$ and ${A}_{i}$ are all constant. The new decision variable $x$ consists of all lower triangular elements of $L_{\mathcal{E}}$, $b_{\mathcal{E}}$, $t$ of (V), and the elements added in order to deal with constraint (36b) in form of hypograph of geometric mean. We have $\bar{n}=O(n^{2})$ because an $n$ dimensional ellipsoid already has $n(n+3)/2$ variables. For convenience, we set $t$ to be the $\bar{n}$-th element of $x$, then we have
@@
Figure 3: Illustration of bottom-up strategy. Although the calculation process is bottom-up, we describe it from top to bottom for clarity of presentation. In the figure, our goal is to compute the MVIE of a closed set $\bar{\mathcal{P}}^{5}$ containing $5$ elements. As shown in the gray box, first we calculate the MVIE for each of its subsets $\bar{\mathcal{P}}_{j}^{4}(1\leq j\leq 5)$ containing $4$ elements, and calculate the MENN of $\bar{\mathcal{P}}^{5}$. Then among these ellipses, we select the largest one that satisfies the constraint of being included by $\bar{\mathcal{P}}^{5}$ as the MVIE of $\bar{\mathcal{P}}^{5}$. Except for the MENN of $\bar{\mathcal{P}}^{5}$ which can be directly calculated by (VI-A), the MVIE of $\bar{\mathcal{P}}_{j}^{4}$ requires to be obtained in a similar way as used for the MVIE of $\bar{\mathcal{P}}^{5}$. As shown in the orange bow, a specific procedure for calculating the MVIE of $\bar{\mathcal{P}}_{1}^{4}$ is given, where its subsets are denoted by $\bar{\mathcal{P}}_{k}^{3}(1\leq k\leq 4)$. Similarly, we compute the MVIE for each subsets $\bar{\mathcal{P}}_{k}^{3}$, and the MENN of $\bar{\mathcal{P}}_{1}^{4}$. Then among the $5$ results, we select the largest one that satisfies the constraint of being included by $\bar{\mathcal{P}}_{1}^{4}$ as the MVIE of $\bar{\mathcal{P}}_{1}^{4}$. If we extend the content of the diagram downwards and upwards: 1) Since at least $3$ halfspaces are required to form a closed space in 2-D, the MVIE of $\bar{\mathcal{P}}_{k}^{3}$ can be directly computed by (VI-A), without the need of splitting and enumeration in the grey or orange box. 2) For a set $\mathcal{P}$ containing more than $5$ elements, its MVIE can be obtained by calculating the MVIE of all its subsets $\bar{\mathcal{P}}^{5}$ and selecting the largest ellipse contained within $\mathcal{P}$.
Then we generalize the Affine Scaling (AS) [38, 39], an interior point method, for efficiently solving the SOCP (V). At each iteration, AS uses the logarithmic barrier function of the constraints to compute a strictly feasible region and calculates the optimal step within this region. For (V), its logarithmic barrier function $\phi(x)$ is defined as
@@
For brevity, we denote $\mathcal{K}$ as the feasible region of (V) , and the interior of the feasible region as $\mathcal{K}^{o}$. Now we obtain the region of AS at $k$-th iteration with the solution $x_{k}\in\mathcal{K}^{o}$ as
$\left\{x~{}|~{}(x-x_{k})^{\mathrm{T}}H_{\mathcal{\phi}}(x_{k})(y-x_{k})\leq 1\right\},$ (44)
which is an ellipsoidal region as $H_{\mathcal{\phi}}(x_{k})$ is positive definite. Thus the update step of (V) can be given by
@@
VI Solving 2-D MVIE Analytically and Rapidly via Randomized Algorithm
For 2-D MVIE, we reformulate it into an enumeration of many subproblems which can be solved analytically. Instead of violent enumeration, we implement a randomized algorithm to achieve an efficient lazy enumeration based on a bottom-up strategy. This lazy randomized algorithm is presented in Sec. VI-A and the analytic solution of the subproblems is detailed in Sec. VI-B.
VI-A Randomized Maximal Inscribed Ellipse Algorithm via Bottom-Up Strategy
In contrast to the numerical solution presented in Sec. V, we propose a novel approach for 2-D MVIE of the convex polygon $\mathcal{P}$, based on the subproblem
¯$\bar{\mathcal{E}}(\mathcal{P}^{N})=\max_{\mathcal{E}}~{}$ ${\operatorname*{aera}({\mathcal{E}})},$ (46a)
@@ VI-A Randomized Maximal Inscribed Ellipse Algorithm via Bottom-Up Strategy
$\mathcal{E}\cap h^{o}\neq\varnothing,~{}\forall~{}h\in{\bar{\mathcal{P}}^{N}},$ (46c)
@@
where ${\bar{\mathcal{P}}^{N}}$ is a subset of $\mathcal{P}$ that is the set of the finite halfspaces whose intersection forms the input polygon, $i$ indicates the amount of elements of ${\bar{\mathcal{P}}^{N}}$, and the boundary of a halfspace $h$ is denoted as $h^{o}$ which is a hyperplane. Since an ellipse has only five degrees of freedom, we require $N\leq 5$. We define $\bar{\mathcal{E}}$ as infinity when the space consisting of the elements of ${\bar{\mathcal{P}}^{N}}$ is not closed. When ${\bar{\mathcal{P}}^{N}}$ is closed and has no redundant element, (VI-A) has an analytical solution, and its result $\bar{\mathcal{E}}$ is the Maximal Ellipse tangent to N edges of the N-gon formed by ${\bar{\mathcal{P}}^{N}}$, denoted as MENN, which is detailed in Sec. VI-B. But if there are redundant elements in a closed ${\bar{\mathcal{P}}^{N}}$, we define its $\bar{\mathcal{E}}$ as a point whose aera is 00. Then based on (VI-A), the MVIE of $\mathcal{P}$ defined in (11) is equivalent to
*$\mathcal{E}^{*}=\max_{{\bar{\mathcal{P}}^{N}}\subseteq\mathcal{P}}{\operatorname*{aera}({\bar{\mathcal{E}}})},~{}\mathrm{s.t.}~{}\bar{\mathcal{E}}\subseteq\mathcal{P},$ (47)
which can be solved through enumeration. Inspired by the subalgorithm of GJK [22], we organize the enumeration in an orderly and recursive way by using a bottom-up strategy, about which we illustrate a detailed example in Fig. 3.
As demonstrated in the orange box of Fig. 3, the MVIE computation of ${\bar{\mathcal{P}}^{4}_{1}}$ that forms a convex quadrilateral are performed. If the MVIE of its any subset ${\bar{\mathcal{P}}_{j}^{3}},1\leq j\leq 4$ satisfies the constraint of being included by ${\bar{\mathcal{P}}^{4}_{1}}$, we can simply select the largest one among the MVIEs of its subsets ${\bar{\mathcal{P}}_{j}^{3}}$ which satisfy this constraint, as the MVIE of ${\bar{\mathcal{P}}^{4}_{1}}$. That is, there is no need to explicitly compute the MENN of ${\bar{\mathcal{P}}^{4}_{1}}$ by (VI-A), which is shown on rightmost side of the orange box. Because the MENN of ${\bar{\mathcal{P}}^{4}_{1}}$ is one of the inscribed ellipses of the triangle formed by one of its subset ${\bar{\mathcal{P}}_{3}^{3}}$. When the largest one of these inscribed ellipses, which is the MVIE of ${\bar{\mathcal{P}}_{3}^{3}}$, is contained by ${\bar{\mathcal{P}}^{4}_{1}}$, it indicates that the MENN of ${\bar{\mathcal{P}}^{4}_{1}}$ is less than or equal to the MVIE of one of its subset ${\bar{\mathcal{P}}_{3}^{3}}$ as shown in Fig. 3. Similarly, this lazy approach can be applied to the MVIE computation of ${\bar{\mathcal{P}}^{5}}$ in the grey box of Fig. 3.
In order to implement above lazy thought, which does not require an exhaustive enumeration, we adopt the MSW algorithm [20], an randomized algorithm that gives an expected time bound linear in the number of constraints for solving LP-type problem [40]. However, with many unclosed cases as shown in Fig. 3, MVIE fails to be directly applied to MSW. We therefore reformulate MVIE into a LP-type problem that satisfies the axioms of MSW. Then we present the reformulation and its implementation into the MSW algorithm.
The new formulation is an optimization problem specified by pairs $(\mathcal{P},w)$, $w:2^{\mathcal{P}}\rightarrow\mathcal{W}$ is a value function. We give three important definitions from MSW: for $\mathcal{H}\subseteq\mathcal{P}$, $h\in\mathcal{P}$,
  • $w(\mathcal{H})$ indicates the value of $\mathcal{H}$,
  • halfspace $h$ is violated by $\mathcal{H}$, if $w(\mathcal{H})<w(\mathcal{H}\cup\{h\})$,
  • the basis of $\mathcal{H}$ is the minimal subset of $\mathcal{H}$ with the same value of $\mathcal{H}$,
based on which, two primitive operations is defined: violation test and basis computation. The value domain $\mathcal{W}$ is defined in lexicographical ordering set $(\mathcal{W},<)$. We denote $<$ as the lexicographical ordering on $\mathbb{R}^{d}$, that is, for $\forall x,y\in\mathbb{R}^{d}$, $x<y$, if $x_{1}<y_{1}$, or $x_{1}=y_{1}$ and $x_{2}<y_{2}$, or so on. We extend the ordering to $\mathbb{R}^{n}\cup\{-\infty\}$ by the convention that $-\infty<x,~{}\forall x\in\mathbb{R}^{d}$, which means a unique minimum value. Then the definition of value function $w$ is given by
[…]
where $\bar{w}_{a}(\mathcal{H})$ is the angle value of $\mathcal{H}$ when the $\mathcal{H}$ is unclosed, and $\bar{w}_{s}(\mathcal{H})$ is the size of the MVIE of $\mathcal{H}$ when closed. Here the objective of reformulated problem is to compute the basis of $\mathcal{P}$, and in the process, the MVIE of $\mathcal{P}$ will be determined. Then we give the definition of angle value of $\mathcal{H}$. We define a new set, denoted as $\mathcal{H}_{v}$, which consists of the vectors corresponding to the direction of each halfspace in $\mathcal{H}$ and give
$\mathcal{C}_{\mathcal{H}}=\left\{{v~{}\big{|}~{}\|v\|\leq 1,~{}v=\sum_{\forall v_{i}\in\mathcal{H}_{v}}k_{i}v_{i},~{}k_{i}\in\mathbb{R}_{\geq 0}}\right\},$ (49)
$\bar{w}_{a}(\mathcal{H})=a_{s}(\mathcal{C}_{\mathcal{H}})-a_{p}(\mathcal{C}_{\mathcal{H}}),$ (50)
where $\mathcal{C}_{\mathcal{H}}$ is a circular sector, $a_{s}(\mathcal{C}_{\mathcal{H}})$ denotes its central angle and $a_{p}(\mathcal{C}_{\mathcal{H}})$ denotes the angle of its polar cone as shown in Fig. 4. Such a particular definition is intended to deal with the corner cases that arise due to the parallel elements in $\mathcal{H}$.
Figure 4: Illustration of the definition of angle value of $\mathcal{H}$. The blue circular sector represents $\mathcal{C}_{\mathcal{H}}$, the green region represents the polar cone of $\mathcal{C}_{\mathcal{H}}$. (b)-(d) illustrate the angle value in the special case where $v_{1}$ and $v_{2}$ are parallel but in opposite directions.
The maximum cardinality of any basis is denoted as combinatorial dimension $\delta$, and based on the value function (48) we have $\delta=5$. According to the lemma of MSW, the reformulated problem can be easily transformed into a basis-regular LP-type problem by adding an element to the end of the lexicographical ordering of $w(\mathcal{H})$ when $\mathcal{H}\neq\varnothing$. The element is given by
$\min\{\delta,|\mathcal{H}|\},$ (51)
where $|\mathcal{H}|$ denotes the number of the elements of $\mathcal{H}$. Now the problem is adapted to the MSW algorithm, and thus can be solved with finite primitive operations, whose expected number is linear in $|\mathcal{H}|$.
Input: $\mathcal{H},\mathcal{X},v_{\mathcal{X}}$
Output: $\mathcal{B},v_{\mathcal{B}}$
@@
2       $\mathbf{{UpdateValue}}(\mathcal{X},v_{\mathcal{X}})$ if $|\mathcal{X}|\geq\delta$ then

### v2

@@ V Solving SOCP-Reformulation of MVIE via Affine Scaling Algorithm
In this section, we propose a practically efficient algorithm for MVIE in a low dimension but with a large number of constraint, which is required in FIRI (Line 1 of Algorithm 1). By carefully handling the orthogonal constraints on the elliptical matrix and the objective function involving the matrix determinant, we reformulate the MVIE problem into a Second-Order Conic Programming (SOCP) form. Then we employ Affine Scaling [45] algorithm for solving this SOCP, achieving efficient solution of MVIE.
V-A SOCP-Reformulation of MVIE
@@
where $\operatorname*{diag}({\cdot})$ indicates either constructing a diagonal matrix or taking all diagonal entries of a square matrix. However, this program still imposes orthonormality constraints on $A_{\mathcal{E}}$. It is worth mentioning that if we use $A_{\mathcal{E}}D_{\mathcal{E}}$ as the decision variable, (27) becomes an SDP with a challenging orthogonal constraint, which is often used to solve MVIE [1, 20, 23].
Then we eliminate the orthogonal constraints (27d). Noting that $A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}$ is always positive definite for the optimal solution of the non-degenerate problem (27). Thus the Cholesky factorization $A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}=L_{\mathcal{E}}L_{\mathcal{E}}^{\mathrm{T}}$ is unique [47] for this solution, where $L_{\mathcal{E}}$ is a lower triangular matrix with positive diagonal entries. If we treat $L_{\mathcal{E}}$ as decision variables, since $A_{\mathcal{E}}$ is orthonormal, the objective (27a) and constraints (27b) can be written as
$\det({D_{\mathcal{E}}})=\sqrt{\det({A_{\mathcal{E}}D_{\mathcal{E}}^{2}A_{\mathcal{E}}^{\mathrm{T}}})}=\det({L_{\mathcal{E}}}),$ (28)
$[{[{A_{\mathcal{P}}A_{\mathcal{E}}D_{\mathcal{E}}}]^{2}\mathbf{1}}]^{\frac{1}{2}}=[{[{A_{\mathcal{P}}L_{\mathcal{E}}}]^{2}\mathbf{1}}]^{\frac{1}{2}}.$ (29)
Consequently, the orthonormality constraint (27d) on $A_{\mathcal{E}}$ is avoided, and (27) is equivalent to
@@
$~{}L_{\mathcal{E}}~{}\mathrm{is~{}lower~{}triangular}.$ (30c)
Now we simplify the objective function $\det({L_{\mathcal{E}}})$ which is the product of the diagonal entries. Denoting the hypograph of geometric mean by
@@
then maximizing the ellpsoid $\mathcal{E}$ is equivalent to maximizing a new variable $t$ with the constraint $({\operatorname*{diag}({L_{\mathcal{E}}}),t})\in\mathcal{K}_{1/n}$. As for the constraints (27b), we denote the SOC as $\mathcal{K}_{n}$ and describe the Cartesian product of $m$ SOC as $\mathcal{K}_{n}^{m}$,
@@
Then the optimization (30) is formulated as
@@
$~{}L_{\mathcal{E}}~{}\mathrm{is~{}lower~{}triangular},$ (34d)
where $m$ denotes the number of halfspaces that compose the convex polytope $\mathcal{P}$ defined in (10). Additionally, $\mathcal{K}_{1/n}$ can also be represented by SOC using additional $O(n)$ variables and cones of $\mathcal{K}_{3}$ [20]. Eventually, we transform (34) into an optimization with only SOC form constraints and a simple linear objective function.
V-B Affine Scaling for Solving SOCP with Massive Constraints
As we reformulate MVIE into a pure SOCP from (34), for brevity, we denote it as
@@ V-B Affine Scaling for Solving SOCP with Massive Constraints
where $x\in\mathbb{R}^{\bar{n}}$, $c_{\mathcal{K}}\in\mathbb{R}^{\bar{n}}$, ${c}_{i}\in\mathbb{R}^{\bar{n}}$, ${d}_{i}\in\mathbb{R}$ and ${A}_{i}$ are all constant. The new decision variable $x$ consists of all lower triangular elements of $L_{\mathcal{E}}$, $b_{\mathcal{E}}$, $t$ of (34), and the elements added in order to deal with constraint (34b) in form of hypograph of geometric mean. We have $\bar{n}=O(n^{2})$ because an $n$ dimensional ellipsoid already has $n(n+3)/2$ variables. For convenience, we set $t$ to be the $\bar{n}$-th element of $x$, then we have
@@
Then we aim to efficiently solve the SOCP (35). Although the dimension $n$ of the original problem (11) is limited to $2$ or $3$, the large constraint number $m$ results in a huge number of constraints $\bar{m}$ in the newly constructed SOCP. To tackle this multi-constraint SOCP, we generalize Affine Scaling (AS) [45], an interior-point method originally used for LP, to SOCP. This algorithm has an closed-form update step of each iteration and exhibits superlinear convergence [48]. At each iteration, AS uses the logarithmic barrier function of the constraints to compute a strictly feasible region and calculates the optimal update step within this region. For (35), its logarithmic barrier function $\phi(x)$ is defined as
@@
Now we obtain the strictly feasible region of AS at $(k+1)$-th iteration based on the feasible $k$-th solution $x_{k}$ as
$\left\{x\in\mathbb{R}^{n}~{}|~{}(x-x_{k})^{\mathrm{T}}H_{\mathcal{\phi}}(x_{k})(y-x_{k})\leq 1\right\},$ (42)
which is an ellipsoidal region as $H_{\mathcal{\phi}}(x_{k})$ is positive definite. Then the update step of (35) can be given by
@@
VI Solving 2-D MVIE Analytically via A Linear-time Complexity Algorithm
Inspired by the linear-time algorithm of the dual problem Minimum Volume Enclosing Ellipsoid (MVEE) [24], in this section, we focus on the construction of an algorithm that achieves linear-time complexity for 2-D MVIE, leveraging its special LP-type problem [25] structure. We begin by providing the necessary background knowledge related to LP-type problems and analyzing the limitations of existing general solutions for applying to MVIE in Sec. VI-A. Then we address these limitations and propose our algorithm with linear time complexity in Sec. VI-B. Finally we provide the analytic solution of the subproblems required in the proposed algorithm in Sec. VI-C.
VI-A Background of LP-type Problem
Let us consider an abstract minimization [25] specified by pairs $(\mathcal{H},w)$, where $\mathcal{H}$ is a finite set of constraints and $w:2^{\mathcal{H}}\rightarrow\mathcal{W}$ is a value function that maps subsets of $\mathcal{H}$ to values in a ordered set $(\mathcal{W},<)$, which has a unique minimum value $\infty$- ∞. For the sake of simplicity in subsequent descriptions, we define finite sets $\mathcal{G}$, $\mathcal{F}$ and a constraint $h$, which satisfy $\mathcal{G}\subseteq\mathcal{F}\subseteq\mathcal{H}$, $h\in\mathcal{H}$. The problem with $(\mathcal{H},w)$ can be considered as an LP-type problem as long as the following two properties are satisfied [25]: for any $\mathcal{G},\mathcal{F}$ and $h$ we have
[…]
  • Locality: with $-\infty<w(\mathcal{G})=w(\mathcal{F})$, if $w(\mathcal{F})<w(\mathcal{F}\cup\{h\})$, then $w(\mathcal{G})<w(\mathcal{G}\cup\{h\})$.
Three important definitions of LP-type problem are given:
  • $w(\mathcal{H})$ is called the value of $\mathcal{H}$,
  • constraint $h$ is violated by $\mathcal{H}$, if $w(\mathcal{H})<w(\mathcal{H}\cup\{h\})$,
  • the basis of $\mathcal{H}$ is the minimal subset of $\mathcal{H}$ with the same value of $\mathcal{H}$.
Then two primitive operations is defined: given a basis $\mathcal{B}$,
  • Violation test: determine whether $w(\mathcal{B})<w(\mathcal{B}\cup\{h\})$, for a constraint $h\notin\mathcal{B}$,
  • Basis computation: compute the basis of $\mathcal{B}\cup\{h\}$ when $h$ is violated by the basis $\mathcal{B}$.
To give readers an intuitive understanding, we provide a specific example: for the dual problem MVEE [24], $\mathcal{H}$ is the input points set and $w(\mathcal{H})$ is the area of the minimum ellipse that can contain all the input points.
The goal of LP-type problem is to compute the basis of the input set $\mathcal{H}$ and its corresponding value $w(\mathcal{H})$. A generalized algorithm for the LP-type problem is given by Matoušek et al. [26]. The algorithm operates by randomly selecting a constraint $h$ from the input set $\mathcal{H}$ and a known basis $\mathcal{B}\subseteq\mathcal{H}$ and performing a violation test. If a violation is detected, a basis of $\mathcal{B}\cup\{h\}$ is computed, then the algorithm is recursively called with the new basis and the set of checked constraints. This algorithm provides an efficient solution to LP-type problems with finite primitive operations, whose expected number is linear in the input set number $|\mathcal{H}|$, thanks to its randomized recursive structure. We refer the reader interested in this complexity conclusion to Matoušek’s work [26].
However, for 2-D MVIE, the aforementioned generalized algorithm described above is not available for two reasons: i) In MVIE, the subset of input constraints may not form a closed polygon, which leads to an undefined ellipse. Using the volume of the ellipse, similar to its dual problem MVEE, to define the value function $w(\cdot)$ is not feasible. ii) There is currently no known method for performing analytical basis computation directly for any subsets in MVIE.
Figure 4: Illustration of a specific instance of the bottom-up strategy. The large orange box specifies the process described by (49) for calculating the MVIE of $\bar{\mathcal{P}}^{4}_{1}$ based on the MVIE of its subsets $\bar{\mathcal{P}}^{3}_{j},1\leq j\leq 4$ and the MENN of itself. The red ellipses indicate MVIE $\mathcal{E}^{*}$ or MENN $\bar{\mathcal{E}}$, depending on the equation beneath each subset. Using the subset in the lower left corner as an example, the solid blue lines indicate elements in the subset $\bar{\mathcal{P}}^{3}_{1}$, the black dashed lines indicate elements in $\bar{\mathcal{P}}^{4}_{1}$ that are not included in the subset $\bar{\mathcal{P}}^{3}_{1}$. Similarly, the blue box illustrates the process of calculating the MVIE of $\bar{\mathcal{P}}^{5}$.
VI-B Randomized Maximal Inscribed Ellipse Algorithm
In this subsection, we address the aforementioned challenges, and then we present a linear-time complexity algorithm by improving the generalized algorithm [26] for 2-D MVIE.
VI-B1 Value Function
We design an additional evaluation of the proximity to closure for the unclosed case and use the MVIE volume for the closed case to construct a new value function that satisfies the properties of the LP-type problem. We define the value domain $\mathcal{W}$ in lexicographical ordering set $(\mathcal{W},<)$. We denote $<$ as the lexicographical ordering on $\mathbb{R}^{2}$, that is, for $\forall x=(x_{1},x_{2})^{\mathrm{T}},y=(y_{1},y_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$, $x<y$, if $x_{1}<y_{1}$, or $x_{1}=y_{1}$ and $x_{2}<y_{2}$. This ordering is extended to $\mathbb{R}^{2}\cup\{-\infty\}$ by the definition that $\infty$- ∞ is the unique minimum value. Considering the frequent failure to form closed polygon for the subsets of constraints in MVIE, we define the value function as
[…]
where $\bar{w}_{u}(\mathcal{G})$ represents a measure of the proximity to closure when $\mathcal{G}$ is not closed. To define $\bar{w}_{u}(\mathcal{G})$, we first define a new set $\mathcal{G}_{n}$ that does not contain duplicate elements, and all elements in the new set $\mathcal{G}_{n}$ come from the unit normal vectors corresponding to the halfspace constrain elements in the set $\mathcal{G}$. Then we define $\bar{w}_{a}(\mathcal{G})$ in lexicographical ordering by
$\bar{w}_{u}(\mathcal{G})=(-a_{p}(\mathcal{G}_{n}),~{}\min\{|\mathcal{G}_{n}|,3\}),$ (45)
where $a_{p}(\mathcal{G}_{n})$ denotes the angle of polar cone of $\mathcal{G}_{n}$. Such a particular definition is intended to deal with the corner cases that arise due to the parallel elements in $\mathcal{G}$.
As for $\bar{w}_{c}(\mathcal{G})$ when $\mathcal{G}$ is closed, we define it by
$\bar{w}_{c}(\mathcal{G})=(\operatorname*{area}(\mathcal{E}^{*}(\mathcal{G}))^{-1},~{}\xi),$ (46)
whose first element represents the inverse of the MVIE’s area of $\mathcal{G}$, and the second element $\xi$ serves to extend this function to $\mathbb{R}^{2}$. The first element of $\bar{w}_{c}(\mathcal{G})$ is always greater than 0, and the first element of $\bar{w}_{u}(\mathcal{G})$ is always less than or equal to 0. Thus no matter what value $\xi$ takes, we have $\bar{w}_{c}>\bar{w}_{u}$. Now based on this value function that satisfies monotonicity and locality, whose proof is detailed in Appendix -A and -B, MVIE can be treated as an LP-type problem.
VI-B2 Basis Computation
To perform analytical basis computation for MVIE, we decompose MVIE into several subproblems that can be solved analytically, then MVIE of $\mathcal{P}$ defined in (11) is equivalent to
$\mathcal{E}^{*}(\mathcal{P})=\max_{{\bar{\mathcal{P}}^{N}}\subseteq\mathcal{P}}{\operatorname*{area}\left({\bar{\mathcal{E}}(\bar{\mathcal{P}}^{N})}\right)},~{}\mathrm{s.t.}~{}\bar{\mathcal{E}}\subseteq\mathcal{P},$ (47)
where $\mathcal{P}$ is the input polygon and also denotes the set of the finite halfspaces whose intersection forms the input polygon in this section. ${\bar{\mathcal{P}}^{N}}$ is its subset containing $N$ elements. Additionally, ¯$\bar{\mathcal{E}}(\bar{\mathcal{P}}^{N})$ is the subproblem defined by
¯$\bar{\mathcal{E}}(\bar{\mathcal{P}}^{N})=\max_{\mathcal{E}}~{}$ ${\operatorname*{area}({\mathcal{E}})},$ (48a)
@@ VI-B2 Basis Computation
$\mathcal{E}~{}\mathrm{is~{}tangent~{}to}~{}h,~{}\forall~{}h\in{\bar{\mathcal{P}}^{N}}.$ (48c)
@@
Since at least $3$ halfspaces are required to form a closed space in 2-D and an ellipse has only $5$ degrees of freedom, we require $3\leq N\leq 5$. When ${\bar{\mathcal{P}}^{N}}$ is closed and has no redundant element, (48) has an analytical solution, and its result $\bar{\mathcal{E}}$ is the Maximal Ellipse tangent to N edges of the N-gon formed by ${\bar{\mathcal{P}}^{N}}$, denoted as MENN, which is detailed in Sec. VI-C. Note that $\mathcal{E}^{*}$ and $\bar{\mathcal{E}}$ are used to denote MVIE and MENN, respectively. For the two cases where the subproblems (48) are infeasible, we handle them with special treatment: i) If ${\bar{\mathcal{P}}^{N}}$ is not closed, we define its $\bar{\mathcal{E}}$ as $\infty$. ii) If there are redundant elements in a closed ${\bar{\mathcal{P}}^{N}}$, that is, these elements are not any side of the formed closed polygon. At this point, it is impossible for these elements to satisfy (48c), then we also define its $\bar{\mathcal{E}}$ as $\infty$. Based on the subproblem (48), MVIE (47) can be solved through enumerating subproblems (48) which has analytical solution.
Inspired by the distance algorithm of GJK [27], we organize the enumeration in an orderly way by using a bottom-up strategy, about which we illustrate a detailed example in Fig. 4. In Fig. 4, we use subscript to distinguish between different subsets of a set, for instance, ${\bar{\mathcal{P}}^{4}_{1}}$ denotes one of the subsets of ${\bar{\mathcal{P}}^{5}}$. Specifically regarding the bottom-up strategy, for a subset ${\bar{\mathcal{P}}^{4}_{1}}$ containing $4$ elements, we compute its MVIE by
$\mathcal{E}^{*}(\bar{\mathcal{P}}^{4}_{1})=\max_{\mathcal{E}}~{}$ ${\operatorname*{area}({{\mathcal{E}}})},$ (49a)
$\mathrm{s.t.}~{}~{}$ ${\mathcal{E}}\subseteq\bar{\mathcal{P}}^{4}_{1},$ (49b)
{¯${\mathcal{E}}\in\{\bar{\mathcal{E}}(\bar{\mathcal{P}}^{4}_{1})\}\cup\{\mathcal{E}^{*}(\bar{\mathcal{P}}^{3}_{j})\},1\leq j\leq 4,$ (49c)
where $\bar{\mathcal{P}}^{3}_{j}$ denotes the subset, which contains $3$ elements, of $\bar{\mathcal{P}}^{4}_{1}$. This equation (49) corresponds to the process in the large orange box in Fig. 4: among the MVIE for each subset $\bar{\mathcal{P}}_{j}^{3}$, and the MENN of $\bar{\mathcal{P}}^{4}_{1}$, we select the largest one that satisfies the constraint (49b) as the MVIE of $\bar{\mathcal{P}}_{1}^{4}$. Based on the result we can build upwards MVIE of $\mathcal{P}^{5}$ containing $5$ elements in a similar way, as illustrated in the blue box of Fig. 4. Eventually, for the input set $\mathcal{P}$ containing more than $5$ elements, its MVIE can be obtained by calculating the MVIE of all its subsets $\bar{\mathcal{P}}^{5}$ and selecting the largest ellipse contained within $\mathcal{P}$. Additionally, it should be noted that the MENN of a closed $\bar{\mathcal{P}}^{3}_{j}$ is its MVIE  [49]. To summarize, we can now compute the MVIE of the input set $\mathcal{P}$ based entirely on its subsets’ MENN ($3\leq N\leq 5$) which can be solved analytically in an efficient enumeration order.
Notion: combinatorial dimension $\delta=5$
Input:   $\mathcal{H}$: input set,  $\mathcal{X}$: a subset of the basis of $\mathcal{H}$
Output: $\mathcal{B}$: basis of $\mathcal{H}$,  $v_{\mathcal{B}}$: value of $\mathcal{B}$
@@
2       $v_{\mathcal{X}}\leftarrow w(\mathcal{X})$   /* defined in (44) */
3       if $|\mathcal{X}|\geq\delta$ then